OpenStudy (anonymous):

Can someone please check my answer. What are the x and y intercepts of the graph 5x-4y=-20

pooja195 (pooja195):

You need to replace x and y with 0

OpenStudy (anonymous):

y- intercept 5x-4y=-20 5x-4y(0)=-20 5x=-20 x=-4 x-intercept 5(0)-4y=-20 -4y=-20 y=5 My answer is (-4,5)

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

@Vocaloid

pooja195 (pooja195):

5x-4(0)=-20 <---solve for x 5(0)-4y=-20 <---solve for y

pooja195 (pooja195):

Ok yes your answer is correct :)

OpenStudy (anonymous):

Ohh okay thanks lol

pooja195 (pooja195):

:P yw